OpenStudy (anonymous):

Solve the equation x+1/3= 2 for x. Please show your work and solution in the space provided. Okay either I need to go back to school or something cuz my little sisters homework is hard.

hero (hero):

x + 1/3 = 2 x = 2 - 1/3 x = 6/3 - 1/3 x = 5/3

hero (hero):

This is hard? Are you sure?

OpenStudy (anonymous):

Im not that smart. I only start college next august.

hero (hero):

Well, you better start hitting the books real fast

OpenStudy (anonymous):

Yeah no kidder right?